OpenStudy (anonymous):

A steam catapult launches a jet aircraft (from rest) from the aircraft carrier John C. Stennis, giving it a speed of 200 mi/h in 3.00 s. (a) Find the average acceleration of the plane.

OpenStudy (anonymous):

assuming acceleration is constant, find the distance the plane moves.

OpenStudy (anonymous):

average acceleration = velocity/ time I believe it would be 200/3 = 66.67 mi/h^2 But I was never good with physics, so I'm not sure.

OpenStudy (anonymous):

And isn't distance = rate * time? d = (200)*3 = 600 miles, since acceleration is constant.